OpenStudy (anonymous):

Solve. 4y + 16 > –8 A. y > –6 B. y > –2 C. y > 2 D. y > 6

OpenStudy (anonymous):

4y+16>-8 So 4y>-8-16 4y>-24 Thus y=-24/4 y=-6

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

You are welcome.